OpenStudy (anonymous):

In this problem, we study the motion of a pendulum that slows down! This is called damped harmonic motion. You start by pulling the pendulum to an angle of 1 radian and dropping it. The length of the pendulum is l = 10 meters. After 1 period, the pendulum swings to a maximum of 0.8 radians. The pendulum’s motion is modeled by y = A cos(t).The amplitude is a decreasing function modeled by an exponential: A = C e^(r)(t) Using the given information, find the values of C and r ! go!

OpenStudy (dumbcow):

Ok by amplitude I am assuming the height the pendulum reaches in the vertical direction. angle is zero means amplitude is zero Ok initial values. t=0, A = C From the fact the angle is 57.3 degrees...i prefer degrees. the height from the top of pendulum is 10*cos(57.3) = 5.4 10 - 5.4 = 4.6 thus A = C = 4.6 Now 1 period implies t = 2pi the angle is now 45.84 degrees the height from the top of pendulum is 10*cos(45.84) = 6.97 10 - 6.97 = 3.03 A = 4.6*e^(2pi*r) = 3.03 Solve for r e^(2pi*r) = .659 2pi*r = -.417 r = -.066 Now we have equation for A \[A = 4.6e^{-.066t}\]